The demand and supply equations are Qd = 100 – 10P and Qs = 30 + 5P. Compute for the


Price equilibrium and quantity equilibrium.

1
Expert's answer
2021-10-31T18:27:55-0400

"Q_D=Q_S\\100-10P=30+P\\100-30=P+10P\\70=11P\\P=\frac{70}{11}\\P=6.37"


"Q_D=100-10(6.363)=36.37\\Q_S=30+6.37=36.37"



"P_e=6.37\\Q_e=36.37"